OpenStudy (anonymous):

consider f(x)= 9-e^x Find the slope of the graph of f(x) at the point where the graph crosses the x-axis. 1-slope=? Find the equation of the tangent line to the curve at this point. 2-y=? Find the equation of the line perpendicular to the tangent line at this point. (This is the normal line.) 3-y=?

OpenStudy (anonymous):

the graph crosses the x-axis at the point e^x=9 x=ln9 1- in order to find the slope take the derivative: f'(x) = -e^x now plug in the ln9 : f'(ln9) = -9 2-tangent line well we have the slope which is -9 and we have the point : (ln9,0) now the general equation for a line is : y = y1 + m(x-x1) plug in the values: y = -9x +9ln9 3- perpendicular line will be opposite and reciprocal to the slope of the tangent line so m(perpendicular) = 1/9 note ( m(perpendicular) * m(tangent) = -1 ) now as we did in 2 : y = x/9 - ln9/9